# Smaller and More Flexible Cuckoo Filters

### v1

@@ Abstract
Cuckoo filters are space-efficient approximate set membership data structures with a controllable false positive rate (FPR) and zero false negatives, similar to Bloom filters. In contrast to Bloom filters, Cuckoo filters store multi-bit fingerprints of keys in a hash table using variants of Cuckoo hashing, allowing each fingerprint to be stored at a small number of possible locations. Existing Cuckoo filters use fingerprints of $(k+3)$ bits per key and an additional space overhead factor of at least $1.05$ to achieve an FPR of $2^{-k}$. For $k=10$, this amounts to $1.365\,kn$ bits to store $n$ keys, which is better than $1.443\,kn$ bits for Bloom filters. The $+3$ for the fingerprint size is required to balance out the multiplied FPR caused by looking for the fingerprint at several locations. In the original Cuckoo filter, the number of hash table buckets is restricted to a power of 2, which may lead to much larger space overheads, up to $2.1\,(1+3/k)\,kn$ bits.
We present two improvements of Cuckoo filters. First, we remove the restriction that the number of buckets must be a power of 2 by using a different placement strategy. Second, we reduce the space overhead factor of Cuckoo filters to $1.06\,(1+2/k)$ by using overlapping windows instead of disjoint buckets to maintain the load threshold of the hash table, while reducing the number of alternative slots where any fingerprint may be found.
A detailed evaluation demonstrates that the alternative memory layout based on overlapping windows decreases the size of Cuckoo filters not only in theory, but also in practice. A comparison with other state-of-the art filter types, Prefix filters and Vector Quotient filters (VQFs), shows that the reduced space overhead makes windowed Cuckoo filters the smallest filters supporting online insertions, with similarly fast queries, but longer insertion times. The improved Cuckoo filters are also more flexible in the sense that Prefix filters and VQFs only support a restricted set of FPRs.
@@ 1 Introduction
Probabilistic filters are space efficient data structures for fast set membership queries. They have many practical applications, such as for database systems [24], networks [14, 25], storage systems [21] and sequence analysis in computational biology [17, 20, 23]. A probabilistic filter represents a set $K$ of keys from a key universe $\mathcal{U}$ and supports at least two operations: (1) inserting a key, and (2) querying if a key is contained in the filter. Instead of storing an exact representation of the set $K$, the keys are fuzzily represented in a way that queries for keys that were inserted into the filter are always correctly answered (no false negative results). Queries about keys that were not inserted, are correctly answered most of the time, but sometimes incorrectly positively. The probability of obtaining a false positive answer is called the false positive rate (FPR). The FPR is controlled by the user and related to the space requirements of the filter. To represent an arbitrary set of keys $K$ with cardinality $|K|=n$ with an FPR of $\varepsilon=2^{-k}$, the theoretically optimal space requirement is $nk$ bits. Practical implementations of filters need $Cnk$ bits with an overhead factor $C>1$ that differs between filter types. Different filter types try to achieve small $C$ and fast insertion and query times; often with different trade-offs between space and time efficiency. In addition, some filter variations support additional operations, such as deletion of keys (dynamic filters), counting the number of times a key was inserted (counting filters), or even storing arbitrary values with the keys (probabilistic key-value stores). Other variants optimize the space requirements for a fixed set of keys, where the complete set of keys has to be known before filter construction starts (static filters).
Filter variants.
However, alternative filter designs often perform better than Bloom filters. Cuckoo filters [11], (Vector) Quotient filters [13, 28] and Prefix filters [10] store a $k$-bit fingerprint in a hash table instead of distributed bits in a bit array. This design allows for storing additional values together with the key fingerprints, simply by using additional bits, giving us not only a probabilistic set membership data structure, but also a probabilistic key-value store without additional effort. The FPR is controlled by the fingerprint size, where larger fingerprints lead to smaller FPRs. Different filter types in this class differ in their hash collision resolution strategies. For example, Cuckoo filters resolve collisions using Cuckoo hashing [27, 11] and Vector Quotient filters use ideas from Robin Hood hashing [28].
The above filters, even the non-static ones, require knowledge of the cardinality $n=|K|$ (or an accurate estimate of it) prior to filter construction. In contrast, extendable filters, such as Infini filters [4], Aleph filters [5], dynamic Cuckoo filters [3] or consistent Cuckoo filters [22] increase their capacity if more keys arrive than originally planned for, at the cost of a higher memory usage and/or higher FPR. Adaptive Cuckoo filters [26] can react to false positive queries by moving them to a hash table. This leads to a lower FPR for future queries at the cost of a larger memory overhead.
@@
Table 1: Overview of non-extendable filter types supporting online insertions. The overhead factor $C=C(k)>1$ specifies the required space per key; one needs $Cnk$ bits to store $n$ keys with an FPR of $2^{-k}$. For some filters, $C(k)$ can vary between the given formula (best case) and twice that value (worst case; here given as worst-case $C(8)$ for a typical $k=8$ scenario). The number of cache misses for insertion/lookup is a proxy for insertion/lookup time.
@@
Standard Cuckoo [11] $1.050\,(1+3/k)$ (†) 2.887 1 to $M$ (∗) 2
Improved Cuckoo [this] $1.057\,(1+2/k)$ 1.321 1 to $M$ (∗) 2
Quotient [1] $1.053\,(1+2.125/k)$ (†) 2.665 some (∗∗) some (∗∗)
Vector Quotient [28] $1.075\,(1+2.914/k)$ (†) 2.933 2 2
@@ 2.1 Multi-way Cuckoo hashing
In standard Cuckoo hashing [27], a set of keys $K=\{x_{1},x_{2},\dots,x_{n}\}\subset\mathcal{U}$ is stored (exactly) in a hash table with $s\geq n$ slots; indexed $10,1,2,\dots,s-$ 1. After inserting $n$ keys, the fill rate (or load factor) of the hash table is $r:=n/s$. The hash table may store additional data associated with each key. Each key $x$ may be inserted into one of two possible slots, computed using two hash functions $f_{1}:\mathcal{U}\to[s]:=\{0,\dots,s-1\}$ and $f_{2}:\mathcal{U}\to[s]$, randomly chosen from a universal family. If both slots $f_{1}(x)$, $f_{2}(x)$ for key $x$ are already occupied, one of the keys at $f_{1}(x)$ or $f_{2}(x)$ is removed and re-inserted into its alternative slot. This slot might also be occupied, so a chain of removals and re-insertions starts until a free slot is found. This can lead to long walks or end in a cycle, and the maximum load factor that can be achieved with a high probability in this setting is $1/2$, a rather low value.
@@
    $(d,l)$ windowed Cuckoo hashing works similarly to $(d,l)$ bucketed Cuckoo hashing, but uses overlapping windows instead of non-overlapping buckets [19]. Each key can be inserted into one of $dl$ slots, but each window overlaps by $l-1$ positions with its surrounding windows, hence the total number of windows is $W=s-l+1$.
@@ 2.2 Cuckoo filters
The two bucket addresses are constrained such that one can be computed from the other and the fingerprint. In the standard Cuckoo filter [11], a hash function $f_{1}:\mathcal{U}\to[B]$ computes the first bucket address of a key $x$, given by $b_{1}=f_{1}(x)$. The second bucket address is the bit-wise XOR of a hash of the fingerprint $h(f_{0}(x))$ and the first bucket address $b_{1}$, where $h:[2^{q}]\to[B]$. Conversely, the first bucket address is obtained back by XORing $b_{2}$ with the same fingerprint hash. The XOR operation is only guaranteed to return a valid address if the number of buckets is a power of 2. A proposed simplification uses the $q$-bit fingerprint $f_{0}(x)$ directly instead of a hash value $h(f_{0}(x))$ [9].
@@
Inserting a key $x$ works as follows: If any of the $2l$ slots where the fingerprint of $x$ can be stored is empty, it is stored there, giving priority to slots in the first bucket $b_{1}$. If all $2l$ slots are full, pick a random one of these slots, remove the stored fingerprint, say of key $x^{\prime}$, and insert fingerprint $f_{0}(x)$ at the now free slot. Then, attempt to insert the removed fingerprint $f_{0}(x^{\prime})$ into its alternative bucket (which can be computed from its former address and the fingerprint itself, without knowledge of $x^{\prime}$). This procedure is repeated for up to a given number of steps. The insertion fails if no free slot can be found. The probability of failure approaches zero if the number of buckets is large enough (and enough steps are allowed). In the described $(2,4)$ configuration, $B=1.02\cdot n/4$ buckets, or $1.02\,n$ slots are sufficient in theory. In practice, in order to limit the length of the random insertion walks, the number of buckets and slots is chosen larger, and the original work [11] uses $1.05\,n$ slots. It has to be noted that a Cuckoo filter (if not over-provisioned) cannot tolerate many additional insertions; these will simply fail.
@@ 3.1 Moving between alternative buckets or windows
We use a different approach, consisting of the first bucket hash function $f_{1}:\mathcal{U}\to[B]$ and an offset hash function $f_{2}:[2^{q}]\to[B-1]$, mapping any fingerprint to an offset less than $B$. For a key $x$, we have the first bucket address $b=f_{1}(x)$ and $b^{\prime}:=(b+1+f_{2}(f_{0}(x)))\bmod B$ with $b^{\prime}\neq b$. From $b^{\prime}$, bucket $b$ cannot be reconstructed symmetrically, but asymmetrically as $b=(b^{\prime}-1-f_{2}(f_{0}(x)))\bmod B$. Hence, we need to use a bit in addition to the fingerprint to store the current bucket choice ($b$ or $b^{\prime}$) that tells us whether we need to add or subtract the offset to obtain the alternative bucket.
It seems disadvantageous that we need one extra bit in each slot. However, when we query the presence of a key, we only consider the key to be present if both the fingerprint and the choice bit are identical. Therefore, we can use the choice bit as one of the 3 extra bits needed to counteract the 8 possible slots. We do not require that this bit is uniformly distributed between 0 and 1 across the filter: Assume that fraction $p$ of all fingerprints (of length $q=k+2$) are stored in their respective first bucket with choice bit 0, and the remaining fraction $1-p$ in their second bucket with choice bit 1. A false positive occurs if we find the fingerprint together with choice bit 0 in the first bucket (probability $\leq 4\cdot 2^{-q}\cdot p$) or with choice bit 1 in the second bucket (probability $\leq 4\cdot 2^{-q}\cdot(1-p)$). So the total probability is bounded by $4\cdot 2^{-q}\cdot(p+1-p)=2^{-k}$, independently of $p$.
@@ 3.2 Windowed layout
The standard Cuckoo filter uses (2,4) bucketed Cuckoo hashing [11]. Alternatively, the array may be divided into $W$ overlapping windows of size $l$, where each window overlaps with the previous window by $l-1$ slots. Using windows has been shown to increase the load threshold for Cuckoo hash tables in comparison to buckets [31, 19]; see Table 2. This allows us to move from $(2,4)$ bucketed Cuckoo hashing to $(2,2)$ windowed Cuckoo hashing without sacrificing much of the load threshold.
@@
Inserting or querying a key works analogously to the version with buckets. To insert a key, check if any of the $2l$ slots is empty. If yes, insert the fingerprint into the empty slot. Otherwise, remove a random fingerprint and try to re-insert it into an alternative slot. To re-insert the removed fingerprint, compute both its current and alternative window based on the stored bits (Figure 2). Run this loop of removal and re-insertion until an empty slot is found, but at most for a fixed number of steps.
For queries, return True if and only if any of the $2l$ possible slots contain the correct fingerprint and the correct choice and window offset bits for the currently searched slot.
@@
    two slots in windows achieve a load threshold ($0.965$) close to four slots in buckets ($0.980$), so we can reduce the number of offsets bits from 2 to 1, saving space. More precisely, filling the hash tables slightly below their theoretical load thresholds, the overhead factor is reduced (for typical not too large $k$) from $1.05\cdot(1+3/k)$ in [11] to $1.06\cdot(1+2/k)$, e.g. from $1.365$ to $1.272$ for $k=10$ (an FPR of $1/1024$).
@@ 4 Implementation Details
We here discuss our implementation using optimized just-in-time compiled Python, how we distinguish empty from full slots in the hash table and how we execute bit-parallel queries over several slots. Further aspects, the parallelization over independent subfilters and the concretely used hash functions, are discussed in Appendix A and B, respectively.

### v2

@@ Abstract
Cuckoo filters are space-efficient approximate set membership data structures with a controllable false positive rate (FPR) and no false negatives, similar to Bloom filters. In contrast to Bloom filters, Cuckoo filters store multi-bit fingerprints of keys in a hash table using variants of Cuckoo hashing, allowing each fingerprint to be stored at a small number of possible locations. Existing Cuckoo filters use fingerprints of $(k+3)$ bits per key and an additional space overhead factor of at least $1.05$ to achieve an FPR of $2^{-k}$. For $k=10$, this amounts to 1.365kn1.365\,kn bits to store $n$ keys, which is better than 1.443kn1.443\,kn bits for Bloom filters. The +3+3 for the fingerprint size is required to balance out the multiplied FPR caused by searching for the fingerprint at $2^{3}=8$ locations. In the original Cuckoo filter, the number of hash table buckets is restricted to a power of 2, which may lead to much larger space overheads, up to 2.1(1+3/k)kn2.1\,(1+3/k)\,kn bits.
We present two improvements of Cuckoo filters. First, we remove the restriction that the number of buckets must be a power of 2 by using a different placement strategy. Second, we reduce the space overhead factor of Cuckoo filters to 1.06(1+2/k)1.06\,(1+2/k) by using overlapping windows instead of disjoint buckets to maintain the load threshold of the hash table, while reducing the number of alternative slots where any fingerprint may be found. The is parallelized implementation using a producer-consumer method which uses lock-free communication via message buffers.
A detailed evaluation demonstrates that the alternative memory layout based on overlapping windows decreases the size of Cuckoo filters not only in theory, but also in practice. A comparison with other state-of-the art filter types, Prefix filters and Vector Quotient filters (VQFs), shows that the reduced space overhead makes windowed Cuckoo filters the smallest filters supporting online insertions, with similarly fast queries, but longer insertion times. The improved Cuckoo filters are also more flexible in the sense that current implementations of Prefix filters and VQFs only support a restricted set of FPRs.
@@ 1 Introduction
are space efficient data structures for approximate fast set membership queries. They have many practical applications, such as for database systems [25], networks [15, 26], storage systems [22] and sequence analysis in computational biology [18, 21, 24]. A probabilistic filter represents a set $K$ of keys from a key universe $\mathcal{U}$ and supports at least two operations: (1) inserting a key, and (2) querying if a key is contained in the set. Instead of storing an exact representation of the set $K$, the keys are fuzzily represented in a way that queries for keys that were inserted into the filter are always correctly answered (no false negative results). Queries about keys that were not inserted are correctly answered most of the time, but sometimes incorrectly positively. The probability of obtaining a false positive answer is called the false positive rate (FPR). The FPR is controlled by the user and related to the space requirements of the filter. To represent an arbitrary set of keys $K$ with cardinality $|K|=n$ with an FPR of $\varepsilon=2^{-k}$, the theoretically optimal space requirement is $nk$ bits. Practical implementations of filters need $Cnk$ bits with an overhead factor $C>1$ that differs between filter types. Different filter types try to achieve small $C$ and fast insertion and query times; often with different trade-offs between space and time efficiency. In addition, some filter variations support additional operations, such as deletion of keys (dynamic filters), counting the number of times a key was inserted (counting filters), or even storing arbitrary values with the keys (probabilistic key-value stores). Other variants optimize the space requirements for a fixed set of keys, where the complete set of keys has to be known before filter construction starts (static filters).
@@
Table 1: Overview of non-extendable filter types supporting online insertions. The overhead factor $C=C(k)>1$ specifies the required space per key; one needs CnkCnk bits to store $n$ keys with an FPR of $2^{-k}$. If a filter only supports a number of slots that is a power of $2$, $C(k)$ can vary between the given formula (best case) and twice that value (worst case; here given as worst-case $C(8)$ for a typical $k=8$ scenario). The number of cache misses for insertion/lookup is a proxy for insertion/lookup time.
@@
Standard Cuckoo [12] 1.050(1+3/k)1.050\,(1+3/k) (†) 2.887 1 to $M$ (∗) 2
Improved Cuckoo [this] 1.057(1+2/k)1.057\,(1+2/k) 1.321 1 to $M$ (∗) 2
Quotient [1] 1.053(1+2.125/k)1.053\,(1+2.125/k) (†) 2.665 some (∗∗) some (∗∗)
Vector Quotient [29] 1.075(1+2.914/k)1.075\,(1+2.914/k) (†) 2.933 2 2
@@
Filter variants.
However, alternative filter designs often perform better than Bloom filters. Cuckoo filters [12], Morton filters [3], (Vector) Quotient filters [14, 29] and Prefix filters [11] store a $k$-bit fingerprint in a hash table instead of distributed bits in a bit array. This design allows for storing additional values together with the key fingerprints, simply by using additional bits, giving us not only a probabilistic set membership data structure, but also a probabilistic key-value store without additional effort. The FPR is controlled by the fingerprint size, where larger fingerprints lead to smaller FPRs. Different filter types in this class differ in their hash collision resolution strategies. For example, Cuckoo filters resolve collisions using Cuckoo hashing [28, 12] and Vector Quotient filters use ideas from Robin Hood hashing [29]. Depending on the collision resolution strategy, the filters achieve different load thresholds. For a hash table with $s$ slots, the load (or fill rate) $r$ is given by $r:=n/s$, where $n$ is the number of inserted keys. The load threshold (or maximum load factor) is the largest load $r$ for which the filter can be successfully constructed with high probability. To achieve low space overhead we need to design a data structure that has a high load threshold while at the same time limiting the number of possible positions a key’s fingerprint may be stored, as each alternative position would increase the FPR if not compensated for by additional bits per key.
The above filters, even the non-static ones, require knowledge of the cardinality $n=|K|$ (or an accurate estimate of it) prior to filter construction. In contrast, extendable filters, such as Infini filters [5], Aleph filters [6], dynamic Cuckoo filters [4] or consistent Cuckoo filters [23] increase their capacity if more keys arrive than originally planned for, at the cost of a higher memory usage and/or higher FPR. Adaptive Cuckoo filters [27] can react to false positive queries by re-arranging fingerprints in the Cuckoo filter. To recognize a false positive, adaptive Cuckoo filters require the whole key set to be additionally stored in a hash table, and are thus only practical in applications where the Cuckoo filter only acts as a fast pre-filter.
@@ 2.1 Multi-way Cuckoo hashing
In standard Cuckoo hashing [28], a set of keys $K=\{x_{1},x_{2},\dots,x_{n}\}\subset\mathcal{U}$ is stored (exactly) in a hash table with $s\geq n$ slots; indexed $0,1,2,\dots,s-1$. After inserting $n$ keys, the fill rate (or load factor) of the hash table is $r:=n/s$. The hash table may store additional data associated with each key. Each key $x$ may be inserted into one of two possible slots, computed using two hash functions $f_{1}:\mathcal{U}\to[s]:=\{0,\dots,s-1\}$ and $f_{2}:\mathcal{U}\to[s]$, randomly chosen from a universal family. If both slots $f_{1}(x)$, $f_{2}(x)$ for key $x$ are already occupied, one of the keys at $f_{1}(x)$ or $f_{2}(x)$ is removed and re-inserted into its alternative slot. This slot might also be occupied, so a chain of removals and re-insertions starts until a free slot is found. This can lead to long walks or end in a cycle, and the maximum load factor that can be achieved with a high probability in this setting is $1/2$, a rather low value.
@@
    $(d,l)$ windowed Cuckoo hashing works similarly to $(d,l)$ bucketed Cuckoo hashing, but uses overlapping windows instead of non-overlapping buckets [20]. Each key can be inserted into one of $dl$ slots, but each window overlaps by l1l-1 positions with its surrounding windows, hence the total number of windows is W=sl+1W=s-l+1.
@@ 2.2 Cuckoo filters
Since only store the fingerprint is stored, the two bucket addresses must be restricted so that one can be calculated from the other and the fingerprint. If we do not want to sacrifice a bit that remembers whether the fingerprint is stored in its first or second bucket, we must use a symmetric function to compute both buckets. The standard Cuckoo filter solves this problem as follows [12]. A hash function $f_{1}:\mathcal{U}\to[B]$ computes the first bucket address of a key $x$, given by $b_{1}=f_{1}(x)$. The second bucket address is the bit-wise XOR of a hash of the fingerprint $h(f_{0}(x))$ and the first bucket address $b_{1}$, where $h:[2^{q}]\to[B]$. Conversely, the first bucket address is obtained back by XORing $b_{2}$ with the same fingerprint hash. The XOR operation is only guaranteed to return a valid address if the number of buckets is a power of 2. A proposed simplification [10] uses the $q$-bit fingerprint $f_{0}(x)$ directly instead of a hash value $h(f_{0}(x))$.
@@
Inserting a key $x$ with 2 hash functions and a bucket size of $l$ works as follows. If any of the $2l$ slots where the fingerprint of $x$ can be stored is empty, it is stored there, giving priority to slots in the first bucket $b_{1}$. If all $2l$ slots are full, pick a random one of these slots, remove the stored fingerprint, say of key $x^{\prime}$, and insert fingerprint $f_{0}(x)$ at the now free slot. Then, attempt to insert the removed fingerprint $f_{0}(x^{\prime})$ into its alternative bucket (which can be computed from its former address and the fingerprint itself, without knowledge of $x^{\prime}$). This procedure is repeated for up to a given number of steps. The insertion fails if no free slot can be found. In this case, the filter size must be increased and all insertions repeated. The probability of failure approaches zero if the number of buckets is large enough (and enough steps are allowed). In the described $(2,4)$ configuration, $B=1.02\cdot n/4$ buckets, or 1.02n1.02\,n slots are sufficient in theory to insert all $n$ elements. In practice, in order to limit the length of the random insertion walks, the number of buckets and slots is chosen larger, and the original work [12] uses 1.05n1.05\,n slots. It has to be noted that a Cuckoo filter (if not over-provisioned) cannot tolerate many additional insertions; these will simply fail.
@@ 3.1 Moving between alternative buckets or windows
We use a different approach, consisting of the first bucket hash function $f_{1}:\mathcal{U}\to[B]$ and an offset hash function $f_{2}:[2^{q}]\to[B-1]$, mapping any fingerprint to an offset less than $B$. For a key $x$, we have the first bucket address $b=f_{1}(x)$ and $b^{\prime}:=(b+f_{2}(f_{0}(x)))\bmod B$ with $b^{\prime}\neq b$. From $b^{\prime}$, bucket $b$ cannot be reconstructed symmetrically, but asymmetrically as $b=(b^{\prime}-f_{2}(f_{0}(x)))\bmod B$. Hence, we need to use a bit in addition to the fingerprint to store the current bucket choice ($b$ or $b^{\prime}$) that tells us whether we need to add or subtract the offset to obtain the alternative bucket.
It seems disadvantageous that we need one extra bit in each slot. However, when we query the presence of a key, we only consider the key to be present if both the fingerprint and the choice bit are identical. Therefore, we can use the choice bit as one of the 3 extra bits needed to counteract the 8 possible slots. We do not require that this bit is uniformly distributed between 0 and 1 across the filter: Assume that fraction $p$ of all fingerprints (of length $q=k+2$) are stored in their respective first bucket with choice bit 0, and the remaining fraction 1p1-p in their second bucket with choice bit 1. A false positive occurs if we find the fingerprint together with choice bit 0 in the first bucket (probability $\leq 4\cdot 2^{-q}\cdot p$) or with choice bit 1 in the second bucket (probability $\leq 4\cdot 2^{-q}\cdot(1-p)$). So the total probability is bounded by $4\cdot 2^{-q}\cdot(p+1-p)=2^{-k}$, independently of $p$.
@@ 3.2 Windowed layout
The standard Cuckoo filter uses (2,4) bucketed Cuckoo hashing [12]. Alternatively, the array may be divided into $W$ overlapping windows of size $l$, where each window overlaps with the previous window by l1l-1 slots. Using windows has been shown to increase the load threshold for Cuckoo hash tables in comparison to buckets [32, 20]; see Table 2. This allows us to move from $(2,4)$ bucketed Cuckoo hashing to $(2,2)$ windowed Cuckoo hashing while maintaining a comparable load threshold.
@@
Inserting or querying a key works analogously to the version with buckets. To insert a key, check if any of the $2l$ slots is empty. If yes, insert the fingerprint into the empty slot. Otherwise, remove a random fingerprint and try to re-insert it into an alternative slot. To re-insert the removed fingerprint, compute both its current and alternative window based on the stored bits (Figure 2). Run this loop of removal and re-insertion until an empty slot is found, but at most for a fixed number of steps. For queries, return True if and only if any of the $2l$ possible slots contain the correct fingerprint and the correct choice and window offset bits for the currently searched slot.
@@
    we reduce the number of offset bits from 2 to 1, saving space, because the load threshold for Cuckoo hashing with two slots in windows and for four slots in buckets are comparable ($0.965$ vs. $0.980$). More precisely, filling the hash tables slightly below their theoretical load thresholds, the overhead factor is reduced from $1.05(1+3/k)$ in [12] to $1.06(1+2/k)$, e.g., from $1.365$ to $1.272$ for $k=10$ (an FPR of $1/1024$).
@@ 4 Implementation Details
We now discuss our implementation using optimized just-in-time compiled Python, how we distinguish empty from full slots in the hash table, how we execute bit-parallel queries over several slots, and how we parallelize insertion. Further aspects, such as the concretely used hash functions, are discussed in Appendix A.